# Preconditioning FEM discretisations of the high-frequency Maxwell equations by either perturbing the coefficients or adding absorption

### v1

@@ 4 Recap of properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$
Let ${\mathcal{A}}_{\ell}:{\cal H}\to{\cal H}^{*}$ be the operator associated to $a_{\ell}(\cdot,\cdot)$; i.e., $\langle{\mathcal{A}}_{\ell}u,v\rangle_{{\cal H}^{*}\times{\cal H}}=a_{\ell}(u,v)$ for all $u,v\in{\cal H}$. We recall the following standard result.
@@
See, e.g., [22, Lemma 6.5.3], [33, Theorem 2.1.44]. The first bound in (4.1) is that $\|{\mathcal{A}}_{\ell}u\|_{{\cal H}^{*}}\geq\gamma^{-1}\|u\|_{{\cal H}}$ for all $u\in{\cal H}$, and the second bound in (4.1) implies that the kernel of ${\mathcal{A}}_{\ell}^{*}$ is empty, and thus the image of ${\mathcal{A}}_{\ell}$ equals ${\cal H}^{*}$; see, e.g., [28, Theorem 2.13(iii)]. ∎
@@
When $\mu_{\ell}$ and $\epsilon_{\ell}$ are both constant multiples of the identity in part of the domain, $\|{\mathcal{A}}^{-1}_{\ell}\|\geq Ck$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [1, Lemma 3.10] (for Helmholtz) and [4, §1.4.1], [30, Example 3.4] (for Maxwell).
For the scattering problem, if the scatterer is nontrapping then $\|{\mathcal{A}}^{-1}_{\ell}\|\leq Ck$; this has been proved in a wide variety of Helmholtz cases (see the literature review in [19]) and for the Maxwell problem with certain nontrapping coefficients in [4] or a nontrapping PEC obstacle in [36, §2]. Under trapping, $\|{\mathcal{A}}^{-1}_{\ell}\|\gg k$; see the results and literature reviews in [25], [2].
For the PML problem, [13, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz radial PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem – indeed, this result was recently proved for the case of no scatterer and Cartesian PML in [6, Lemma 10].
@@ 5.1 Notation for the Galerkin matrices
Let the matrix $\mathsf{D}$ be such that
$\left\|v_{N}\right\|_{{\cal H}}^{2}=\left\|\mathbf{V}\right\|_{\mathsf{D}}^{2}\quad\text{ for }v_{N}=\sum_{j=1}^{N}V_{j}\phi_{j}\,\text{ with }\,V_{j}\in\mathbb{C}^{N}.$ (5.3)
We also use the notation $\|\cdot\|_{\mathsf{D}}$ to denote the induced norm on matrices. Finally, let $\|\cdot\|_{2}$ denote the Euclidean norm on $\mathbb{C}^{N}$, and let $m_{\pm}$ be such that
@@
i.e., the quantity $(m_{+}/m_{-})^{2}$ (whose square root appears in the bound (5.8) below) is the ratio of the largest to the smallest eigenvalue – i.e., the condition number – of the mass matrix $\mathsf{M}$.
@@
    §5.3 gives sufficient conditions for (5.5a) to hold (including a new result on this); i.e., conditions under which the discrete inf-sup constant is bounded by the continuous inf-sup constant.
@@
Remark 5.4 (Recovering the Helmholtz results of [16] about preconditioning with absorption).
The set up in [16] considers preconditioning the Helmholtz equation with absorption by letting $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$. In Corollary 5.3 this corresponds to setting $\alpha:=\varepsilon/k^{2}$. By Remark 4.3, $\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\sim k$ when the Helmholtz problem is nontrapping. The right hand side of (5.10) is then proportional to $\varepsilon/k$, which is the result proved in [16, Theorem 1.4].
@@
    (as a consequence of the Elman estimate [8, 7]; see [16, Corollary 1.9]/[20, Corollary 5.5]).
@@
We highlight that, at least in the case of preconditioning by absorption, the action of $\mathsf{A}_{2}^{-1}$ must be further approximated to provide a practical preconditioner (as discussed in §1).
@@
(see Remark 6.3). For the Helmholtz case with $\mu_{1}=\mu_{2}$, [20, Lemma 2.7] exhibits $\epsilon_{1},\epsilon_{2}$ such that the bound (5.14) is sharp in its $k$-dependence (in this example $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}\sim k$).
@@
By using standard finite-element inverse estimates, [20, Lemma 5.3] proved (for the Helmholtz case) bounds on $\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}}$, $\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}}$, $\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}$, and $\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{2}$ with $L^{q}$ norms of both $\mu^{-1}_{1}-\mu^{-1}_{2}$ and $\epsilon_{1}-\epsilon_{2}$ appearing on the right-hand sides. These bounds also hold for the Helmholtz and Maxwell problems in §3; however, we do not include them here since (i) these bounds involves powers of $h^{-1}$ (from the inverse estimates) and are thus worse than (5.7) and (5.8), and (ii) the bounds (5.7) and (5.8) are sufficient for studying preconditioning with absorption (becoming (5.9) and (5.10) in Corollary 5.3) and, in particular, verifying the assumption about this in [26].
@@
(i) Theorem 5.2 covers all the Helmholtz problems in §3.1, whereas [16] and [20] only consider the exterior Dirichlet problem with a nontrapping obstacle (with the obstacle allowed to be empty). For this problem, [16] considers truncation of the exterior domain by an impedance boundary condition, and [20] considers truncation by either an impedance boundary condition or the exact Dirichlet-to-Neumann map.
(ii) Theorem 5.2 makes assumptions only about $a_{1}(\cdot,\cdot)$ (i.e., the problem we want to precondition), whereas [16] and [20] make assumptions about both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$; e.g., the mesh threshold in [16, Equation 1.12] is for the problem with absorption, i.e., $a_{2}(\cdot,\cdot)$, and [20, Theorems 2.2 and 2.3] make assumptions on both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$.
(iii) Theorem 5.2 is proved under the discrete inf-sup condition (5.5) (which, by Theorem 4.1, is equivalent to the Galerkin matrix being invertible), whereas the bounds of [16] and [20] are proved under stronger assumptions than a discrete inf-sup condition. Indeed, [16, Equations 1.13 and 1.14] prove bounds analogous to (5.8) under the assumption that, given any data and a sequence of finite-dimensional subspaces, the sequence of Galerkin solutions is quasi-optimal, with constant independent of $k$ – we see in Corollary 5.10 below that this implies that the Galerkin matrix is invertible, with the norm of its inverse bounded by the norm of the PDE solution operator. Furthermore, [20, Lemma 5.3] proves bounds analogous to (5.7) and (5.8) under assumptions weaker than quasi-optimality, but stronger than invertibility of the Galerkin matrix; see [20, Condition 4.2].
5.3 Conditions under which the discrete inf-sup constant is proportional to the (continuous) inf-sup constant
@@
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big{|}a_{\ell}(u_{h},v_{h})\big{|}}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\big{(}1+C_{\rm G2,\ell}C_{\rm sol,\ell}\big{)}}.$
Several analyses of the $h$-FEM for Helmholtz and Maxwell prove the bound (5.15) directly; e.g., this is done in [27, Theorem 4.1] for the Maxwell impedance problem in §3.2.2 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small.
@@
Suppose that given $k_{0}>0$ there exists $c$ such that $\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\geq c$ for all $k\geq k_{0}$. Given $p\in\mathbb{Z}^{+}$, let $({\cal H}_{h})_{h>0}$ be a sequence of finite-dimensional subspaces of ${\cal H}$.
@@
Since the proofs of Corollaries 5.10 and 5.11 are so short, we give them here.
@@
$a(u,v)=a(\overline{v},\overline{u})\quad\text{ for all }u,v\in{\cal H},$ (5.20)
then the definitions of ${\mathcal{A}}^{*}_{\ell}$ and ${\mathcal{A}}_{\ell}$ imply that $({\mathcal{A}}^{*}_{\ell})^{-1}f=\overline{{\mathcal{A}}_{\ell}^{-1}\overline{f}}$, and thus the bound (5.18) is equivalent to the corresponding bound with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$. The condition (5.20) holds for all the Helmholtz and Maxwell problems in §3 either when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are Hermitian or when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ corresponding to a radial PML (see [14, Lemma 2.3]).
Furthermore, the proof of [15, Theorem 1.7], which establishes (5.18) with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ for general Helmholtz problems, also proves (5.18).
@@
The result that the discrete inf-sup constant is bounded by the continuous inf-sup constant when $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small is [29, Theorem 4.2]. By the Schatz argument [34], [29, Theorem 4.3], if $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small, then the sequence of Galerkin solutions is quasi-optimal, and so [29, Theorem 4.2] is morally equivalent to Corollary 5.10.
@@
  • All the Helmholtz problems in §3 satisfy the assumptions of Corollary 5.11 by [15, Theorems 1.5 and 1.7] (with [15, Theorem 1.5] giving (5.17) and [15, Theorem 1.7] giving (5.18)).
  • For the Maxwell PEC problem of §3.2.1 discretised using Nédélec edge elements of either first or second type, the bound (5.17) under the condition (5.16) is proved in [3, Theorem 1.3]. The bound (5.18) is not proved in [3] (the arguments of [3] use a duality argument that “builds in” the splitting used to prove (5.18)). Therefore, although we expect that the bound (5.15) holds in the preasymptotic regime for the Maxwell PEC problem (via a result analogous to Corollary 5.11), right now [3, Theorem 1.3] only allows us to use Corollary 5.10 in the asymptotic regime (i.e. under the condition that $(kh)^{p}\|{\mathcal{A}}^{-1}_{\ell}\|$ is sufficiently small).
  • As noted above, for the Maxwell impedance problem of §3.2.2, the bound (5.15) is proved directly in [27, Theorem 4.1] for the Maxwell impedance problem in §3 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small (note that $\|{\mathcal{A}}^{-1}_{\ell}\|\sim k$ here by [23]).
@@ 6 Proofs of the main results
Recall from §2 that if $a_{\ell}(\cdot,\cdot)$ satisfies the assumptions in §2, then so does $a^{*}_{\ell}(\cdot,\cdot)$ defined by $a^{*}_{\ell}(u,v)=\overline{a_{\ell}(v,u)}$, with the same constants $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$. Furthermore, if the discrete inf-sup condition (5.5) on $a(\cdot,\cdot)$ holds, then it also holds for $a^{*}(\cdot,\cdot)$, with the same $C_{{\rm dis},N,2}$; see, e.g., [22, Lemma 6.5.3], [33, Remark 2.1.45] (this is just the result that the norm of the adjoint operator equals the norm of the original operator).
@@
We now prove the bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ in (5.7) and (5.8). By Theorem 4.1, the discrete inf-sup condition (5.5) implies that $\mathsf{A}_{2}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ is invertible (see also, e.g., [33, Theorem 4.2.1]). Then, by the definitions of $\mathsf{A}_{\ell}$ (5.1), $\mathsf{S}_{\mu}$, and $\mathsf{M}_{\epsilon}$ (5.2),
@@
$\big{\|}\mathsf{A}_{2}^{-1}\mathsf{S}_{\mu^{-1}}\big{\|}_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\quad\text{ and }\quad\big{\|}\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big{\|}_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$
@@
6.2 Norm of inverse solution operator under perturbation
Lemma 6.2 (Norm of inverse solution operator under perturbation).

### v2

@@ 3 Recap of properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$
In this section, we recap properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$, since it appears in the statement of the main result (Theorem 4.2 below).
[…]
@@
[…]
@@
When $\mu_{\ell}$ and $\epsilon_{\ell}$ are both constant multiples of the identity in part of the domain, $\|{\mathcal{A}}^{-1}_{\ell}\|\geq Ck$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [2, Lemma 3.10] (for Helmholtz) and [5, §1.4.1], [33, Example 3.4] (for Maxwell).
For the scattering problem, if the scatterer is nontrapping then $\|{\mathcal{A}}^{-1}_{\ell}\|\leq Ck$; this has been proved in a wide variety of Helmholtz cases (see, e.g.,, the literature review in [21]) and for the Maxwell problem with certain nontrapping coefficients in [5] or a nontrapping PEC obstacle in [40, §2]. Under trapping, $\|{\mathcal{A}}^{-1}_{\ell}\|\gg k$; see the results and literature reviews in [28], [3].
For the PML problem, [15, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz radial PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem. Indeed, this result was recently proved – up to a factor of $k$ loss – for the case of no scatterer and Cartesian PML in [8, Lemma 10].
@@ 4 Statement of the main results
This section states precisely the main results of the paper; the proofs of these results are then given in §5.
@@ 4.1 Notation for the Galerkin matrices
Let the matrix $\mathsf{D}$ be such that the norm relation (1.3) holds. We also use the notation $\|\cdot\|_{\mathsf{D}}$ to denote the induced norm on matrices. Finally, let $\|\cdot\|_{2}$ denote the Euclidean norm on $\mathbb{C}^{N}$, and let $m_{\pm}$ be such that
@@
i.e., the quantity $(m_{+}/m_{-})^{2}$ (whose square root appears in the bound (4.7) below) is the ratio of the largest to the smallest eigenvalue – i.e., the condition number – of the mass matrix $\mathsf{M}$. For the $h$-FEM with quasi-uniform meshes, the ratio $m_{+}/m_{-}$ is independent of $h$; see, e.g., [18, Equation 4.2], [22, Lemma 5.1], [35, Lemma 4.6].
@@
    §4.3 gives sufficient conditions for (4.4a) to hold (including a new result on this); i.e., conditions under which the discrete inf-sup constant is bounded by a constant times the continuous inf-sup constant.
@@
    (as a consequence of the Elman estimate [10, 9]; see [18, Corollary 1.9]/[22, Corollary 5.5]).
@@
We highlight that, at least in the case of preconditioning by absorption, the action of $\mathsf{A}_{2}^{-1}$ must be further approximated to provide a practical preconditioner (as discussed in §1.1).
@@
(see Remark 5.3). For the Helmholtz case with $\mu_{1}=\mu_{2}$, [22, Lemma 2.7] exhibits $\epsilon_{1},\epsilon_{2}$ such that the bound (4.13) is sharp in its $k$-dependence (in this example $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}\sim k$).
@@
By using standard finite-element inverse estimates, [22, Lemma 5.3] proved (for the Helmholtz case) bounds on $\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}}$, $\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}}$, $\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}$, and $\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2}$ with $L^{q}$ norms of both $\mu^{-1}_{1}-\mu^{-1}_{2}$ and $\epsilon_{1}-\epsilon_{2}$ appearing on the right-hand sides. These bounds also hold for the Helmholtz and Maxwell problems in §2; however, we do not include them here since (i) these bounds involves powers of $h^{-1}$ (from the inverse estimates) and are thus worse than (4.6) and (4.7), and (ii) the bounds (4.6) and (4.7) are sufficient for studying preconditioning with absorption (becoming (4.8) and (4.9) in Corollary 4.3) and, in particular, verifying the assumption about this in [29].
@@
The bounds (4.6)/(4.7) and (4.9) are essentially the same as those proved in [22] and [18], respectively. Nevertheless, Theorem 4.2 improves the results of [18, 22] by enlarging the class of Helmholtz problems to which the bounds apply, and weakening the assumptions on the finite-dimensional space. In more detail:
(i) Theorem 4.2 covers all the Helmholtz problems in Lemma 2.2, whereas [18] and [22] only consider the exterior Dirichlet problem with a nontrapping obstacle (with the obstacle allowed to be empty). For this problem, [18] considers truncation of the exterior domain by an impedance boundary condition, and [22] considers truncation by either an impedance boundary condition or the exact Dirichlet-to-Neumann map.
(ii) Theorem 4.2 makes assumptions only about $a_{1}(\cdot,\cdot)$ (i.e., the problem we want to precondition), whereas [18] and [22] make assumptions about both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$; e.g., the mesh threshold in [18, Equation 1.12] is for the problem with absorption, i.e., $a_{2}(\cdot,\cdot)$, and [22, Theorems 2.2 and 2.3] make assumptions on both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$.
(iii) Theorem 4.2 is proved under the discrete inf-sup condition (4.4) (which, by Theorem 3.1, is equivalent to the Galerkin matrix being invertible), whereas the bounds of [18] and [22] are proved under stronger assumptions than a discrete inf-sup condition. Indeed, [18, Equations 1.13 and 1.14] prove bounds analogous to (4.7) under the assumption that, given any data and a sequence of finite-dimensional subspaces, the sequence of Galerkin solutions is quasi-optimal, with constant independent of $k$ – we see in Corollary 4.9 below that this implies that the Galerkin matrix is invertible, with the norm of its inverse bounded by the norm of the PDE solution operator. Furthermore, [22, Lemma 5.3] proves bounds analogous to (4.6) and (4.7) under assumptions weaker than quasi-optimality, but stronger than invertibility of the Galerkin matrix; see [22, Condition 4.2].
4.3 Auxiliary result: conditions under which the discrete inf-sup constant is proportional to the (continuous) inf-sup constant
@@
[…]
In other words, the bound (4.14) on the discrete solution for the restricted class of data $f\in{\cal H}_{0}$ (instead of $f\in{\cal H}^{*}$) is sufficient to obtain a bound on the discrete inf-sup constant.
Several analyses of the $h$-FEM for Helmholtz and Maxwell prove the bound (4.14) directly; e.g., this is done in [30, Theorem 4.1] for the Maxwell impedance problem in Part (ii) of Lemma 2.4 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small.
@@
[…]
@@
Since the proofs of Corollaries 4.9 and 4.10 are so short, we give them here.
@@
[…]
then the definitions of $\mathcal{A}$$^{*}_{\ell}$ and $\mathcal{A}$$_{\ell}$ imply that $({\mathcal{A}}^{*}_{\ell})^{-1}f=\overline{{\mathcal{A}}_{\ell}^{-1}\overline{f}}$, and thus the bound (4.17) is equivalent to the corresponding bound with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$. The condition (4.19) holds for all the Helmholtz and Maxwell problems in Lemmas 2.2 and 2.4 either when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are Hermitian or when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ corresponding to a radial PML (see [16, Lemma 2.3]).
Furthermore, the proof of [17, Theorem 1.7], which establishes (4.17) with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ for general Helmholtz problems, also proves (4.17).
@@
The result that the discrete inf-sup constant is bounded by the continuous inf-sup constant when $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small is [32, Theorem 4.2]. By the Schatz argument [38], [32, Theorem 4.3], if $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small, then the sequence of Galerkin solutions is quasi-optimal, and so [32, Theorem 4.2] is morally equivalent to Corollary 4.9.
@@
  • All the Helmholtz problems in Lemma 2.2 satisfy the assumptions of Corollary 4.10 by [17, Theorems 1.5 and 1.7] (with [17, Theorem 1.5] giving (4.16) and [17, Theorem 1.7] giving (4.17)).
  • For the Maxwell PEC problem of Part (i) of Lemma 2.4 discretised using Nédélec edge elements of either first or second type, the bound (4.16) under the condition (4.15) is proved in [4, Theorem 1.3]. The bound (4.17) is not proved in [4] (the arguments of [4] use a duality argument that “builds in” the splitting used to prove (4.17)). Therefore, although we expect that the bound (4.14) holds in the preasymptotic regime for the Maxwell PEC problem (via a result analogous to Corollary 4.10), right now [4, Theorem 1.3] only allows us to use Corollary 4.9 in the asymptotic regime (i.e. under the condition that $(kh)^{p}\|{\mathcal{A}}^{-1}_{\ell}\|$ is sufficiently small).
  • As noted above, for the Maxwell impedance problem of Part (ii) of Lemma 2.4, the bound (4.14) is proved directly in [30, Theorem 4.1] for the Maxwell impedance problem in Part (ii) of Lemma 2.4 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small (note that $\|{\mathcal{A}}^{-1}_{\ell}\|\sim k$ here by [26]).
@@ 5 Proofs of the main results
Sections 5.1-5.3 are devoted to the proof of Theorem 4.2, and Section 5.4 is devoted to the proof of Lemma 4.8.
@@
Recall from §2 that if $a_{\ell}(\cdot,\cdot)$ satisfies the assumptions in §2, then so does $a^{*}_{\ell}(\cdot,\cdot)$ defined by $a^{*}_{\ell}(u,v)=\overline{a_{\ell}(v,u)}$, with the same constants $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$. Furthermore, if the discrete inf-sup condition (4.4) on $a(\cdot,\cdot)$ holds, then it also holds for $a^{*}(\cdot,\cdot)$, with the same $C_{{\rm dis},N,2}$; see, e.g., [25, Lemma 6.5.3], [37, Remark 2.1.45] (this is just the result that the norm of the adjoint operator equals the norm of the original operator).
@@
We now prove the bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ in (5.2) and (5.3). By Theorem 3.1, the discrete inf-sup condition (4.4) implies that $\mathsf{A}_{2}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ is invertible. Then, by the definitions of $\mathsf{A}_{\ell}$ (4.1), $\mathsf{S}_{\mu}$, and $\mathsf{M}_{\epsilon}$ (4.2),
@@
$\big\|\mathsf{A}_{2}^{-1}\mathsf{S}_{\mu^{-1}}\big\|_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\quad\text{ and }\quad\big\|\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big\|_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{,}}$
@@
5.2 Norm of solution operator under perturbation
Lemma 5.2 (Norm of solution operator under perturbation).